OpenStudy (anonymous):

Can someone check my work 2. -x^4+7= x^2-2

OpenStudy (anonymous):

\[-x ^{4}+7-(x ^{2}-2)=0\]

OpenStudy (anonymous):

\[-x ^{4}+7-x ^{2}+2\]

OpenStudy (tkhunny):

So far so good, except for the "= 0" that disappeared. It is NOT optional.

OpenStudy (anonymous):

\[-x ^{4}-x ^{2}+9=0\]

OpenStudy (tkhunny):

Well? How about \(x^{4} + x^{2} - 9 = 0\)?

OpenStudy (anonymous):

now im putting it in u^2=x^4 u=x^2 form \[-u ^{2}-u+9\] dividing both sides by -1 \[u ^{2}+u-9\] add 9 to both sides \[u ^{2}+u=9\]

OpenStudy (tkhunny):

That last thing was no good. you need a quadratic equation equal to zero and you just killed it. Go back one step. That's where I left you... \(u^{2} + u - 9 = 0\) Now what?

OpenStudy (anonymous):

adding (1/2)^2 to both sides (this is where i got confused) \[u ^{2}+u+\frac{ 1 }{ 4 }=\frac{ 37 }{ 4 }\] \[(u+\frac{ 1 }{ 4 })^{2}=\frac{ 37 }{ 4 }\] \[u+\frac{ 1 }{ 4 }=\frac{ \sqrt{37} }{ 4 }\] \[u=\frac{ \sqrt{37}-1 }{ 2 }\]

OpenStudy (anonymous):

do i leave it like that or subtract 1 into 37 find the sqrt of 36 and divide by 2?

OpenStudy (tkhunny):

Now, see, you didn't tell me you were "Completing the Square". Fair enough. The quadratic formula is exactly the same thing. third equation is no good. the right-hand side should be \(\dfrac{\sqrt{37}}{2}\). Start again from there.

OpenStudy (tkhunny):

Also, that same one needs a \(\pm\) in there, somewhere.

OpenStudy (anonymous):

can you explain more please why it wouldnt stay 37/4?

OpenStudy (tkhunny):

This is correct: \(\left(u^2 + u +\dfrac{1}{4}\right) = \dfrac{37}{4}\) The next step was not correct. Almost! \(\left(u +\dfrac{1}{2}\right)^{2} = \dfrac{37}{4}\) See that 1/2 in there? You still had 1/4. The next step was a little off. You found the square root fo the numerator, but not the denominator. You also didn't notice that a negative value would suffice. It should be: \(u +\dfrac{1}{2} = \pm\sqrt{\dfrac{37}{4}} = \pm\dfrac{\sqrt{37}}{2}\) Understand? Can you finish?

OpenStudy (anonymous):

im completely lost lol sorry from here would i subtract 1/2 from both sides

OpenStudy (anonymous):

but i do understand why the 4 changed

OpenStudy (tkhunny):

The 4 didn't change. (1/2)^2 = 1/4

OpenStudy (anonymous):

no the \[\sqrt{\frac{ 37 }{ 4 }}\] went to \[\frac{ \sqrt{37} }{ 2 }\] right?

OpenStudy (tkhunny):

\(\sqrt{4} = 2\)

OpenStudy (anonymous):

yes thats what i was getting at now would i subtract 1/2?

OpenStudy (anonymous):

@phi?

OpenStudy (tkhunny):

Why is there a question? You have (u + 1/2). How else will you be solving for 'u'?

OpenStudy (anonymous):

subtracting 1/2

OpenStudy (tkhunny):

Do it. Where does that lead?

OpenStudy (anonymous):

\[u=\frac{ \sqrt{37}-1 }{ 2 }\]

OpenStudy (tkhunny):

No good. You still missed that positive or negative solutions exist for that square root. \(u = \dfrac{-1 \pm \sqrt{37}}{2}\)

OpenStudy (phi):

It looks like you have the right idea, and need to polish up the details

OpenStudy (anonymous):

\[u=\frac{ -1\sqrt{37} }{ 2 } u=\frac{ \sqrt{37}-1 }{ 2 }\]

OpenStudy (tkhunny):

?? That first one doesn't mean anything. Why not write it just like I did? LaTeX \ p m

OpenStudy (phi):

you left out a minus sign in the first solution

OpenStudy (anonymous):

\[u=\frac{ -1-\sqrt{37} }{ 2 }\]

OpenStudy (tkhunny):

It's okay to write the solutions separately, but you probably don't want to do that. If we were solving for "u", we'd be done. We're supposed to be solving for 'x'.

OpenStudy (phi):

to finish, remember that the u = x^2 so \[ x= \pm \sqrt{u} \] you end up with 4 solutions

OpenStudy (anonymous):

this is confusing me

OpenStudy (anonymous):

2 with quadratic formula and 2 with completing the square?

OpenStudy (anonymous):

\[x=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (tkhunny):

Just as a note -- Sometimes, and I'm serious, just remembering what it is you are doing is the hardest part of the problem. We've wandered a long way from where we started. Just go back and review. Get it all in your head, again. It will fit! :-)

OpenStudy (anonymous):

\[x=\frac{ 1\pm \sqrt{-1^{2}-4(-1)(9)} }{ 2(-1) }\]

OpenStudy (anonymous):

\[x=\frac{ 1\pm \sqrt{1+4(9)} }{ -2 }\]

OpenStudy (anonymous):

\[x=\frac{ 1\pm \sqrt{37} }{ -2 }\]

OpenStudy (anonymous):

ty @tkhunny and @phi sorry if im frustrating yall :/

OpenStudy (tkhunny):

Why make it harder than it has to be? Generally, we want to get that leading coefficient positive. -x^4+7= x^2-2 ==> -x^4 - x^2 + 9 = 0 ==> x^4 + x^2 - 9 = 0 \(x^{2} = \dfrac{-1\pm\sqrt{1^{2} - 4(1)(-9)}}{2(1)} = \dfrac{-1\pm\sqrt{37}}{2} \rightarrow x = \pm\sqrt{\dfrac{-1\pm\sqrt{37}}{2} }\) It's not pretty.

OpenStudy (phi):

I don't get frustrated, but sometimes the asker does. you can use either the quadratic formula or complete the square to solve a quadratic. You did that for u^2+u−9=0 but, as you may recall, the question was x^4+x^2−9=0 where you let x^2 = u once you find the values for u, you can solve for the values of x using \[ x= \pm \sqrt{u} \]

OpenStudy (tkhunny):

Personally, I am not a fan of the direct substitution. As you noticed, this enticed you to forget the original problem statement! I MUCH prefer just recognizing the Quadratic and laying it out - as demonstrated above.

OpenStudy (anonymous):

so with completing the square its \[x^2=\frac{ -1\pm \sqrt{37} }{ 2 }\] quadratic \[x=\pm \sqrt{\frac{ -1\pm \sqrt{37} }{ 2 }}\]

OpenStudy (anonymous):

quadratic seems easier than the first one

OpenStudy (anonymous):

\[x=\sqrt{\frac{ -1\pm \sqrt{37} }{ 2 }}\]

OpenStudy (phi):

yes, with a ± in front \[ x=\pm \sqrt{\frac{ -1\pm \sqrt{37} }{ 2 }} \] that is short-hand for 4 separate solutions

OpenStudy (anonymous):

ahh alright i have 1 more question can you help me or walk me through with